OpenStudy (anonymous):

How do I solve this: 4x^2 + 5X =10 please help

OpenStudy (phi):

first add -10 to both sides, to put this in "standard form" \[ 4x^2 + 5x -10 = 10 -10 \] or \[ 4x^2 + 5x -10 =0\] I would now use the quadratic formula See http://www.khanacademy.org/math/trigonometry/polynomial_and_rational/quad_formula_tutorial/v/using-the-quadratic-formula for how to do this